Question Help Assume that different groups of couples use a particular method of gender selection and each couple gives birth to one baby. This method is designed to increase the likelihood that each baby will be a girl, but assume that the method has no effect, so the probability of a girl is 0.5. Assume that the groups consist of 45 couples. Complete parts (a) through (c) below a. Find the mean and the standard deviation for the numbers of girls in groups of 45 births. The value of the mean is μ- Type an integer or a decimal. Do not round.) The value of the standard deviation is (Round to one decimal place as needed.) b. Use the range rule of thumb to find the values separating results that are significantly low or significantly high Values of L1 girls or fewer are significantly low (Round to one decimal place as needed ) Values of□ girls or greater are significantly high Round to one decimal place as needed.) c. Is the result of 36 girls a result that is significanty high? What does it suggest about the effectiveness of the method? The result l ▼| significantly high because 36 girls is suggest that the method (Round to one decimal place as needed.) ▼| □ girls. A result of 36 girls would Part C: The result (is, is not) significantly high, because 36 girls is (greater than, less than, equal to) __ girls. A result of 36 girls would suggest that the method (is not effective, is effective.)
Answer #1

a) \mu = n * p = 45 * 0.5 = 22.5

\sigma = sqrt(n * p * (1 - p)) = sqrt(45 * 0.5 * 0.5) = 3.354

b) Values of girls or fewer significantly low = \mu - 2\sigma = 22.5 - 2 * 3.354 = 15.792

Values of girls or greater are significantly high = \mu + 2\sigma = 22.5 + 2 * 3.354 = 29.208

c) Result is significantly high, because 36 girls is greater than girls. A result of 36 girls would suggest that the method is not effective.